OpenStudy (domebotnos):

1. A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero.

OpenStudy (domebotnos):

My answer choices are 3e2 e 3e None of these

OpenStudy (anonymous):

If \(s(t)\) describes the position, then the derivative of \(s\) gives you the velocity, i.e. \(v(t)=s'(t)\). Find out when the velocity will be zero: \[\begin{align*}v(t)&=s'(t)\\ &=\ln3t+t\frac{3}{3t}\\ &=\ln3t+1 \end{align*}\] Solve for \(t\): \[\ln3t+1=0~~\Rightarrow~~t=\frac{1}{3e}\] The acceleration at time \(t\) is given by the derivative of the velocity function, \(a(t)=v'(t)\): \[a(t)=\frac{3}{3t}=\frac{1}{t}\] Velocity was found to be zero at \(t=\dfrac{1}{3e}\), so the acceleration at this time is \[a\left(\frac{1}{3e}\right)=\frac{1}{\frac{1}{3e}}=3e\]